# Clust-PSI-PFL: A Population Stability Index Approach for Clustered Non-IID Personalized Federated Learning

### v1

@@ I Introduction
Federated learning (FL) [mcmahan2017communication] enables collaborative training across many clients while keeping raw data local by aggregating model updates instead of centralizing examples (so-called CL); however, non-independent and identically distributed (non-IID) data can induce biased updates and degrade global performance [g2024noniiddatafederatedlearning] (see Fig. 1). Personalized FL (PFL) mitigates these issues by tailoring models to clients; methods are commonly grouped into (i) regularization-based approaches that keep local objectives close to the global model, (ii) selection-based schemes that prioritize or filter clients during training, and (iii) clustering-based techniques that partition clients with similar data distributions and train one model per cluster [tan2022towards]. Our approach belongs to the clustering-based family.
Figure 1: (Left) Typical FL training process with FedAvg. (Right) Illustration of weight divergence for FL with non-IID data
This work uses the Population Stability Index (PSI), commonly employed in credit scoring to monitor distributional drift [yurdakul2018statistical, du2023proposed], as a principled tool for client clustering in FL. PSI provides a quantitative measure of cross-client data distributional divergence, which we leverage to identify more homogeneous groups for training. We instantiate and assess this mechanism under label skew, a form of non-IID data known to most strongly degrade FL performance [g2024noniiddatafederatedlearning]. Extending the PSI-based clustering procedure to additional skew types (e.g., attribute or quantity skew) is left for future work.
@@
Diagnosing non-IID data in FL remains an open challenge. Surveys by Pei et al. [pei2024review] and Li et al. [li2020federated] emphasize 4 priorities: (i) lightweight, a priori diagnostics; (ii) privacy-preserving FL analysis; (iii) principled quantification of non-IID degree; and (iv) leveraging these formalizations to design optimization methods with stronger performance and, ideally, convergence guarantees.
@@
To the best of our knowledge, this is the first study to integrate PSI into FL both as a non-IID data quantifier and as a basis for client clustering to enhance global model performance.
@@ II Background and Related Work
In Fig. 1 (left), the standard round-based FL pipeline shows: a server that initializes a global model and broadcasts it to $K$ clients; each trains locally on private data and returns updates (optionally protected via multiparty secure computing (MPC) [bohler2021secure] or differential privacy (DP) [apple_privacy_2017]); the server aggregates the model’s weight (e.g., FedAvg) to form a new global model and rebroadcasts, repeating until convergence.
Non-IID client data is a primary challenge in FL [g2024noniiddatafederatedlearning]. Unlike IID regimes, non-IID data client distributions (Fig. 1, right) induce client drift: local parameters $W_{t}^{i}$ deviate from the global optimum $W_{t}^{(\mathrm{Opt})}$, impeding convergence and yielding non-optimal solutions. In addition, aggregating misaligned updates can drive the global $W_{t}^{(\mathrm{Avg})}$ to converge towards a poor local optimum, further degrading accuracy.
Differences in per-client label distributions are a primary source of difficulty for FL, adversely impacting both accuracy and convergence dynamics [aggregators_FL, hsieh2020non]. To quantify the severity of non-IID data, several measures have been proposed [g2024noniiddatafederatedlearning]. Within PFL, three predominant mitigation families are widely studied: regularization-based techniques, client-selection strategies, and clustering-based solutions [tan2022towards], as discussed in the following sections.
@@
Prior work measures client-distribution mismatch with three common metrics. Hellinger distance (HD) estimates deviation from a balanced reference and has shown strong discrimination, particularly in cross-device FL [tan2023privacy, jimenez2024fedartml]. Jensen–Shannon distance (JSD) supports high-level comparisons and the construction of client-to-client similarity graphs to mitigate label-skew effects [ahmed2023semisupervised, xu2024fblg]. Earth Mover’s/Wasserstein distance (EMD) is used to guide client scheduling and assess non-IID severity by comparing client and global label distributions [chen2022emd].
@@
One of the most widespread solutions to tackle the non-IID data effects is FedProx [li2020federated]. It generalizes FedAvg by adding a proximal term $\mu\geq 0$ that controls the penalization strength. This penalizes drift and stabilizes optimization under non-IID data.
Under the FedOpt framework [reddi2020adaptive], clients perform local stochastic gradient descent (SGD), while the server uses an adaptive update (e.g., Adagrad, Yogi, Adam) to the aggregated weights, thereby improving stability and effectiveness under non-IID data. The variants differ in moment accumulation/normalization (e.g., Yogi’s conservative second moment), offering robustness to sparse gradients and non-convexity and reducing hyperparameter sensitivity compared to FedAvg, as supported by theory and experiments.
FedAvgM [hsu2019measuring] augments FedAvg with server-side momentum, smoothing aggregated updates, and stabilizing training under non-IID data. By maintaining a moving average of the global update direction, it reduces oscillations on skewed partitions and often surpasses FedAvg. The trade-off of FedAvgM is additional hyperparameters (momentum coefficient, server learning rate) that require careful tuning.
@@
To mitigate non-IID effects, Power-of-choice (PoC) [pmlr-v151-jee-cho22a] preferentially samples high-loss clients, balancing convergence speed and bias; it achieves up to $3\times$ faster convergence and $\sim 10\%$ higher test accuracy than random selection. In contrast to PoC, our PSI-guided clustering focuses training within distributionally coherent client clusters, further improving accuracy and efficiency.
HACCS [wolfrath2022haccs] clusters clients by data-histogram similarity and, within each cluster, selects low-latency participants, preserving distributional coverage while favoring fast devices. This yields robustness to individual dropouts (as long as similar clients remain), maintains balanced representation, and accelerates convergence.
FedCLS [li2022fedcls] guides client sampling using group-level label information and Hamming distances between one-hot label vectors, preferentially selecting clients with complementary (diverse) label distributions. Compared to FedAvg with random selection, it achieves faster convergence, higher accuracy, and more stable training (well-suited to non-IID federated settings).
@@
Clustered Federated Learning (CFL), proposed by Sattler et al. [sattler2020clustered], tackles non-IID data by recursively clustering clients via the cosine similarity of their local updates when training stalls (small global norm, large client norm). It bipartitions to minimize cross-cluster similarity and repeats; the method is model-agnostic, FedAvg-compatible, and requires no preset number of clusters.
FedSoft, by Ruan and Joe-Wong [ruan2022fedsoft], extends clustered FL to soft clustering, letting each client be a mixture of source distributions and jointly learning cluster models plus per-client personalized models. It uses a proximal local objective that encodes cluster models for knowledge transfer with per-round cost comparable to standard FL. The method provides convergence guarantees and outperforms hard-clustering baselines on synthetic and real datasets.
@@ III Clust-PSI-PFL Proposed Methodology
This section presents our Clust-PSI-PFL approach; a high-level workflow is shown in Fig. 2. Before training, every client transmits a compact label-frequency count (not raw data) to the server. The server aggregates these frequencies to form a reference label distribution and computes the PSI to quantify each client’s divergence from that reference. Based on the PSI metric, clients are grouped into distributionally homogeneous clusters, and a dedicated model is trained per cluster.
Figure 2: High-level Clust-PSI-PFL training process.
III-A PSI Calculation
In the label-skew setting, the client-wise $PSI$ [siddiqi2012credit] for client $i$ is defined in Eq. 1. This expression coincides with the statistical “Jeffreys divergence” [Jeffreys1946InvariantPrior]:
@@ III-A PSI Calculation
where $P(y=c)$ denotes the global probability mass function (pmf) of label $y$ for class $c$, computed at the server from the aggregated label counts reported by clients; $P_{i}(y=c)$ is the pmf of label $y$ for class $c$ on client $i$; and $C$ is the total number of classes for $y$. The superscript $L$ indicates that the computation pertains to the label-skew setting. As discussed earlier, we focus on label skew; attribute and quantity skew are left for future work.
In addition to the aggregate $PSI_{i}^{L}$, we decompose Eq. 1 into class-wise contributions:
@@
The terms $PSI_{i,c}^{L}$ quantify, for each class $c$, the $i$th client deviation from the global label distribution. We later assemble the client-feature vector $PSI_{i}^{L},\,PSI_{i,1}^{L},\,\ldots,\,PSI_{i,C}^{L}\,]$ for clustering clients.
Since $PSI$ is computed on a per-client basis, the federation-wide degree of non-IID in the label-skew setting is summarized by the weighted average, denoted $WPSI^{L}$, as defined in Eq. 3:
@@
where, $K$ denotes the total number of participating clients, $n_{i}$ is the number of samples on the $i$th client, and $N=\sum_{i=1}^{K}n_{i}$ is the aggregate sample count across all clients. Smaller values of $PSI_{i}$ and $WPSI^{L}$ reflect greater inter-client homogeneity, whereas larger values denote stronger non-IID data.
@@ III-B PSI-based Cluster Creation and Training
Clients are clustered with K-means++, selected for its robust seeding and broad adoption [ArthurVassilvitskii2007KMeansPP]. Each client is encoded as a feature vector that concatenates the overall PSI and the per-class PSI values (Eqs. 1 and 2). These features are standardized to have a mean of zero and a variance of one. This step is lightweight, adding minimal server-side overhead, since its cost scales mainly with the (typically small) number of classes $C$.
@@
We provide a concise analysis of the computational complexity of the Clust-PSI-PFL clustering phase. The cost arises mainly from (i) computing PSI for each client and (ii) selecting the number of clusters via K-means++ with the silhouette score. Computing PSI over $K$ clients and $C$ classes requires $\mathcal{O}(K\,C)$ operations. Let $I$ be the average number of K-means++ iterations, and let the search examine K2K-2 candidate cluster counts $j\in\{2,\ldots,K-1\}$. A single K-means++ run with $j$ clusters costs $\mathcal{O}(I\,K\,j\,(C+1))$, while computing the silhouette costs $\mathcal{O}(K^{2}(C+1))$. Summed over the K2K-2 candidates, the model-selection stage costs $\mathcal{O}\!\big((C+1)\big(I\,K\sum_{j=2}^{K-1}j+(K-2)\,K^{2}\big)\big)=\mathcal{O}\!\big((C+1)\big(I\,K\,((K-1)K/2-1)+(K-2)\,K^{2}\big)\big)$. Therefore, the total clustering complexity is $\mathcal{O}\!\big(K\,C+(C+1)\big(I\,K\,((K-1)K/2-1)+(K-2)\,K^{2}\big)\big)$. In practice, $C$ and the candidate range (K2)(K-2) are small (and silhouettes can be computed on a subsample), so the overhead is modest; subsequent training proceeds as in FedAvg, except that aggregation is performed independently for each of the $\tau$ clusters.
@@
The weighted PSI ($WPSI^{L}$) is well-suited metric to quantify non-IID in FL. We benchmark it against HD, JSD, and EMD [jimenez2024fedartml], standard divergences for comparing distributions (see Section II). Unlike these baselines, $WPSI^{L}$ delivers fine-grained, client-level diagnostics with minimal overhead and, thanks to PSI’s classwise decomposition, enables per-class attribution to pinpoint labels driving divergence, capabilities HD/JSD/EMD generally lack without extra modifications. Thus, $WPSI^{L}$ is an interpretable and efficient primary metric for non-IID assessment.
Figure 3: Relationship between $WPSI^{L}$ and the level of non-IID data for the Dirichlet and Similarity partition protocols.
To evaluate the practical utility of $WPSI^{L}$, we conducted a broad empirical study spanning multiple datasets, levels of non-IID data, partitioning schemes, and random seeds (see Section IV). Centralized datasets were partitioned using the Dirichlet [jimenez2024fedartml] and Similarity [wang2023distribution] partition protocols, after which we computed $WPSI^{L}$, HD, JSD, and EMD. Fig. 3 depicts how these metrics vary with the degree of non-IID data, parameterized by $\alpha$ and $S$. Our findings are as follows: (i) $WPSI^{L}$ exhibits an exponential-decay trend as IID data increases under both protocols; (ii) JSD and HD change more nearly linearly with non-IID data across both schemes; and (iii) EMD is less reliable as a non-IID quantifier due to non-monotonic behavior, yielding identical values at distinct non-IID levels.
Figure 4: Feature importance of models predicting the level of non-IID data for the Dirichlet (top) and Similarity (bottom) partition protocols.

### v2

@@ I Introduction
Federated learning (FL) [22] enables collaborative training across many clients while keeping raw data local by aggregating model updates instead of centralizing examples (so-called CL); however, non-independent and identically distributed (non-IID) data can induce biased updates and degrade global performance [16] (see Fig. 1). Personalized FL (PFL) mitigates these issues by tailoring models to clients; methods are commonly grouped into (i) regularization-based approaches that keep local objectives close to the global model, (ii) selection-based schemes that prioritize or filter clients during training, and (iii) clustering-based techniques that partition clients with similar data distributions and train one model per cluster [30]. Our approach belongs to the clustering-based family.
Figure 1: Standard FedAvg training overview (left) and depiction of model/weight drift under non-IID scenarios (right).
This work uses the Population Stability Index (PSI), commonly employed in credit scoring to monitor distributional drift [36, 8], as a principled tool for client clustering in FL. PSI provides a quantitative measure of cross-client data distributional divergence, which we leverage to identify more homogeneous groups for training. We instantiate and assess this mechanism under label skew, a form of non-IID data known to most strongly degrade FL performance [16]. Extending the PSI-based clustering procedure to additional skew types (e.g., attribute or quantity skew) is left for future work. We deliberately start with label skew, since it is among the most practically impactful and challenging forms of non-IID data in FL. We avoid claiming generality beyond this setting and discuss extensions in Section V.
@@
Diagnosing non-IID data in FL remains an open challenge. Surveys by Pei et al. [23] and Li et al. [21] emphasize 4 priorities: (i) lightweight, a priori diagnostics; (ii) privacy-preserving FL analysis; (iii) principled quantification of non-IID degree; and (iv) leveraging these formalizations to design optimization methods with stronger performance and, ideally, convergence guarantees.
@@
To the best of our knowledge, this is the first study to integrate PSI into FL as a basis for client clustering to enhance global model performance. While PSI has previously been used to guide client selection in non-IID PFL [17], we instead leverage PSI-derived data to cluster clients into distributionally homogeneous groups and train cluster-specific models.
@@ II Background and Related Work
In Fig. 1 (left), the standard round-based FL pipeline shows: a server that initializes a global model and broadcasts it to $K$ clients; each trains locally on private data and returns updates (optionally protected via multiparty secure computing (MPC) [4] or differential privacy (DP) [6]); the server aggregates the model’s weight (e.g., FedAvg) to form a new global model and rebroadcasts, repeating until convergence.
Non-IID client data is a primary challenge in FL [16]. Unlike IID scenarios, non-IID client distributions (Fig. 1, right) induce client drift: starting from the broadcast global model $w_{t}$, local training drives client models $w_{i}^{t+1}$ toward different client-specific optima $w_{i}^{\star}$, rather than toward the global optimum $w^{\star}$. This misalignment makes client updates point in inconsistent directions, slowing convergence and yielding suboptimal solutions. In addition, aggregating these heterogeneous updates can steer the server update away from $w^{\star}$, causing the global iterate (e.g., $w_{t+1}$) to move toward an undesirable region or poor local solution (highlighted in the figure as a red square), thereby further degrading accuracy.
Differences in per-client label distributions are a primary source of difficulty for FL, adversely impacting both accuracy and convergence dynamics [20, 11]. To quantify the severity of non-IID data, several measures have been proposed [16]. Within PFL, three predominant mitigation families are widely studied: regularization-based techniques, client-selection strategies, and clustering-based solutions [30], as discussed in the following sections.
@@
Prior work measures client-distribution mismatch with three common metrics. Hellinger distance (HD) estimates deviation from a balanced reference and has shown strong discrimination, particularly in cross-device FL [31, 15]. Jensen–Shannon distance (JSD) supports high-level comparisons and the construction of client-to-client similarity graphs to mitigate label-skew effects [1, 35]. Earth Mover’s/Wasserstein distance (EMD) is used to guide client scheduling and assess non-IID severity by comparing client and global label distributions [5].
@@
One of the most widespread solutions to tackle the non-IID data effects is FedProx [21]. It generalizes FedAvg by adding a proximal term $\mu\geq 0$ that controls the penalization strength. This penalizes drift and stabilizes optimization under non-IID data.
Under the FedOpt framework [24], clients perform local stochastic gradient descent (SGD), while the server uses an adaptive update (e.g., Adagrad, Yogi, Adam) to the aggregated weights, thereby improving stability and effectiveness under non-IID data. The variants differ in moment accumulation/normalization (e.g., Yogi’s conservative second moment), offering robustness to sparse gradients and non-convexity and reducing hyperparameter sensitivity compared to FedAvg, as supported by theory and experiments.
FedAvgM [12] augments FedAvg with server-side momentum, smoothing aggregated updates, and stabilizing training under non-IID data. By maintaining a moving average of the global update direction, it reduces oscillations on skewed partitions and often surpasses FedAvg. The trade-off of FedAvgM is additional hyperparameters (momentum coefficient, server learning rate) that require careful tuning.
@@
To mitigate non-IID effects, Power-of-choice (PoC) [13] preferentially samples high-loss clients, balancing convergence speed and bias; it achieves up to $3\times$ faster convergence and $\sim 10\%$ higher test accuracy than random selection. In contrast to PoC, our PSI-guided clustering focuses training within distributionally coherent client clusters, further improving accuracy and efficiency.
HACCS [33] clusters clients by data-histogram similarity and, within each cluster, selects low-latency participants, preserving distributional coverage while favoring fast devices. This yields robustness to individual dropouts (as long as similar clients remain), maintains balanced representation, and accelerates convergence.
FedCLS [19] guides client sampling using group-level label information and Hamming distances between one-hot label vectors, preferentially selecting clients with complementary (diverse) label distributions. Compared to FedAvg with random selection, it achieves faster convergence, higher accuracy, and more stable training (well-suited to non-IID federated settings).
@@
Clustered Federated Learning (CFL), proposed by Sattler et al. [27], tackles non-IID data by recursively clustering clients via the cosine similarity of their local updates when training stalls (small global norm, large client norm). It bipartitions to minimize cross-cluster similarity and repeats; the method is model-agnostic, FedAvg-compatible, and requires no preset number of clusters.
FedSoft, by Ruan and Joe-Wong [26], extends clustered FL to soft clustering, letting each client be a mixture of source distributions and jointly learning cluster models plus per-client personalized models. It uses a proximal local objective that encodes cluster models for knowledge transfer with per-round cost comparable to standard FL. The method provides convergence guarantees and outperforms hard-clustering baselines on synthetic and real datasets.
In contrast to CFL/FedSoft, Clust-PSI-PFL clusters clients using label-histogram PSI features computed before any gradient or update is produced, and exploits PSI’s per-class decomposition for interpretable client descriptors. We also select the number of clusters via a silhouette criterion, rather than CFL-style recursive splitting or FedSoft’s soft-mixture optimization.
@@ III Clust-PSI-PFL Proposed Methodology
This section presents our Clust-PSI-PFL approach; a high-level workflow is shown in Fig. 2. Before training, every client transmits a compact label-frequency count (not raw data) to the server. We assume these per-client label counts can be shared in our threat model. When they are deemed sensitive, the same pipeline is compatible with MPC or DP noise added to histogram counts. The server aggregates these frequencies to form a reference label distribution and computes the PSI to quantify each client’s divergence from that reference. Based on the PSI metric, clients are grouped into distributionally homogeneous clusters, and a dedicated model is trained per cluster.
Figure 2: Overview of the Clust-PSI-PFL procedure.
III-A Calculating PSI
In the label-skew case, we quantify how much each client’s label distribution differs from the global one using the client-level $PSI$ [28], given in Eq. 1. This quantity is equivalent to the Jeffreys divergence [14]:
@@ III-A Calculating PSI
Here, P(y=c)P(y{=}c) is the server-side (global) probability mass function (pmf) for class $c$, obtained by aggregating the label counts reported by clients, while $P_{i}(y{=}c)$ is the corresponding pmf on client $i$; $C$ denotes the number of classes. The superscript $L$ emphasizes that we consider label skew (other skew types are left for future work).
To capture which classes contribute most to a client’s mismatch, we also use the per-class terms obtained by isolating the summands in Eq. 1:
@@
The vector $PSI_{i}^{L},\,PSI_{i,1}^{L},\,\ldots,\,PSI_{i,C}^{L}\,]$ is then used as a client descriptor for clustering.
Finally, we summarize federation-level label non-IID data with the sample-size weighted average $WPSI^{L}$:
@@
where $K$ is the number of clients, $n_{i}$ is the local sample count, and $N=\sum_{i=1}^{K}n_{i}$. Smaller $PSI_{i}^{L}$ (and $WPSI^{L}$) indicate more homogeneous clients, whereas larger values reflect stronger non-IID label skew.
@@ III-B PSI-based Cluster Creation and Training
Clients are clustered with K-means++, selected for its robust seeding and broad adoption [2]. Each client is encoded as a feature vector that concatenates the overall PSI and the per-class PSI values (Eqs. 1 and 2). These features are standardized to have a mean of zero and a variance of one. This step is lightweight, adding minimal server-side overhead, since its cost scales mainly with the (typically small) number of classes $C$.
@@
Silhouette score [25]. To quantify the quality of a candidate clustering with $j$ clusters, we use the (average) silhouette score. For each client feature vector $x_{i}$, let $a(i)$ denote the average distance from $x_{i}$ to all other points in its assigned cluster (intra-cluster cohesion), and let $b(i)$ denote the minimum, over all other clusters, of the average distance from $x_{i}$ to the points in that other cluster (nearest-cluster separation). The silhouette coefficient of $x_{i}$ is defined as $\mathrm{sil}(i)=\frac{b(i)-a(i)}{\max\{a(i),b(i)\}}$, which lies in [1,1][-1,1], with larger values indicating better separation and cohesion. We report the mean silhouette over all clients as $s(j)$ and select $\tau=\arg\max_{j}s(j)$.
@@
We provide a concise analysis of the computational complexity of the Clust-PSI-PFL clustering phase. The cost arises mainly from (i) computing PSI for each client and (ii) selecting the number of clusters via K-means++ with the silhouette score. Computing PSI over $K$ clients and $C$ classes requires $\mathcal{O}(K\,C)$ operations. Let $I$ be the average number of K-means++ iterations, and let the search examine K2K-2 candidate cluster counts $j\in\{2,\ldots,K-1\}$. A single K-means++ run with $j$ clusters costs $\mathcal{O}(I\,K\,j\,(C+1))$, while computing the silhouette costs $\mathcal{O}(K^{2}(C+1))$. Summed over the K2K-2 candidates, the model-selection stage costs $\mathcal{O}\!\big((C+1)\big(I\,K\sum_{j=2}^{K-1}j+(K-2)\,K^{2}\big)\big)=\mathcal{O}\!\big((C+1)\big(I\,K\,((K-1)K/2-1)+(K-2)\,K^{2}\big)\big)$. Therefore, the total clustering complexity is $\mathcal{O}\!\big(K\,C+(C+1)\big(I\,K\,((K-1)K/2-1)+(K-2)\,K^{2}\big)\big)$. In practice, the number of clusters explored is small, and this selection is performed only once before FL training, so the added cost is amortized over all communication rounds. In very large-scale deployments, silhouette scores can be approximated on a subsample of clients with negligible impact on the selected number of clusters. Subsequent training proceeds as in FedAvg, except that aggregation is performed independently for each of the $\tau$ clusters.
@@
The weighted PSI ($WPSI^{L}$) is well-suited metric to quantify non-IID in FL. We benchmark it against HD, JSD, and EMD [15], standard divergences for comparing distributions (see Section II). Unlike these baselines, $WPSI^{L}$ delivers fine-grained, client-level diagnostics with minimal overhead and, thanks to PSI’s classwise decomposition, enables per-class attribution to pinpoint labels driving divergence, capabilities HD/JSD/EMD generally lack without extra modifications. Thus, $WPSI^{L}$ is an interpretable and efficient primary metric for non-IID assessment.
Figure 3: Mapping between $WPSI^{L}$ and non-IID level under the Dirichlet and Similarity partition protocols.
To evaluate the practical utility of $WPSI^{L}$, we conducted a broad empirical study spanning multiple datasets, levels of non-IID data, partitioning schemes, and random seeds (see Section IV). Centralized datasets were partitioned using the Dirichlet [15] and Similarity [32] partition protocols, after which we computed $WPSI^{L}$, HD, JSD, and EMD. Fig. 3 shows the relationship between $WPSI^{L}$ and the degree of non-IID data, controlled by the Dirichlet parameter $\alpha$ and the Similarity parameter $S$, with the corresponding values of alternative metrics (HD, JSD, EMD) reported on the secondary axis. Our findings are as follows: (i) $WPSI^{L}$ decreases with increasing $\alpha$ and $S$ (i.e., as data become more IID), following a pronounced decay under both protocols; (ii) JSD and HD vary more smoothly and approximately linearly with $WPSI^{L}$ across both schemes; and (iii) EMD is less reliable as a non-IID quantifier due to non-monotonic behavior in some settings, potentially yielding identical values at distinct non-IID levels.
Figure 4: Feature importances for predicting non-IID level under Dirichlet (top) and Similarity (bottom) partition protocols.